A deep-sea diver uses a gas cylinder with a volume of 10.0 L and a content of 51.1 g of O2 and 33.6 g of He. You may want to reference ( pages 415 - 417) Section 10.6 while completing this problem.

lculate the partial pressure of each gas and the total pressure if the temperature of the gas is 17 ∘C.

Express the pressures in atmospheres to three significant digits separated by commas.

PO2, PHe, Ptotal =
Answer #1

we know that

moles = mass / molar mass

also

molar mass of 02 = 32 g/mol

molar mass of He = 4 g/mol

so

moles of 02 = 51.1 / 32 = 1.596875

moles of He = 33.6 / 4 = 8.4

now

for gases

PV = nRT

given

volume (V) = 10 L

temperature = 17 C = 17 + 273 K = 290 K

R = 0.0821

so using these values

For 02

P02 x 10 = 1.596875 x 0.0821 x 290

P02 = 3.8

so

partial pressure of 02 = 3.8 atm

now

for He

PHe x 10 = 8.4 x 0.0821 x 290

PHe = 19.99956

so

partial pressure of He = 20 atm

now

total pressure = partial pressure of 02 + partial pressure of He

total pressure = 3.8 + 20

total pressure = 23.8

so

total pressure is 23.8 atm